Evaluate 5^(2x+3) = 3^(x-1)
Get bases the same or use ln. The numbers in () are exponents.

Answer by stanbon(57210)

Evaluate 5^(2x+3) = 3^(x-1)
Get bases the same or use ln. The numbers in () are exponents.
-----
Changing the base 5 to base 3:
Let 3^k = 5
k = ln(5)/ln(3)
k = 1.4650
----
Rewrite the problem:
3^[1.4650(2x+3)] = 3^(x-1)
--------------------------------
Equation to solve:
1.4650(2x+3) = x-1
----
2.9299x+ 4.3949 = x-1
----
1.9299x = -5.3949
x = -2.7954
================
Cheers,
Stan H.
=================